OpenStudy (anonymous):

consider the sequence 2,6,18,54... complete each statement, 1) A(1)=2=2 x 3^ 2) A(2)=6=2 x 3= 2 x 3^ 3) A(3)=18=2 x 3 x 3= 2 x 3^ 4) A(4)=54=2 x 3 x 3 x 3= 2 x 3^

OpenStudy (mathmate):

The question is asking you to complete the "number sentence" by filling the missing number at the end of the equation. I'll give you an \(example\) for a similar calculation. 15 = 3 x 5^x divide both sides by 3 soo 5=5^x since 5^1=5, so x=1, or 15=3 x 5^1 Another example: 96=3 x 2^x divide by three, 32=2^x since 2^5 = 32, so 96=3 x 2^5